OpenStudy (anonymous):

An array of pennies has 5 rows and 8 columns. Which is NOT a way to divide the group equally? A. fifths B. eighths C. fourths D. thirds

OpenStudy (anonymous):

D

OpenStudy (zepp):

A way to divide the group equally means that you can't get all row same # of pennies and all columns same # of pennies 5 * 8 = 40 Now you have to find which one doesn't get a quotient belonging to N

OpenStudy (anonymous):

i know is three but, i dont understand that

OpenStudy (zepp):

5 * 8 = 40 A) Fifths; 40/5 = 8 B) Eighths; 40/8 = 5 C) Fourths; 40 / 4 = 10 D) Thirds; 40/3 = 13.3333333333

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks :)